# Enhancing LLM Reliability via Explicit Knowledge Boundary Modeling

### v3

@@ 4.2 Reliability Evaluation
Our approach consistently achieves superior Quality-F1 scores, indicating improved effectiveness in distinguishing predictions as ”sure” or ”unsure” based on confidence levels. Notably, a high indicates reliable ”sure” predictions, fostering user trust with minimal need for verification. Simultaneously, a high ensures that the ”unsure” category encompasses a broad spectrum of potentially correct outputs, forming a solid basis for further refinement. Collectively, these metrics contribute to a higher Optimal-F1 score, representing the theoretical maximum post-refinement performance. In other words, our system optimally balances precision and recall, delivering both immediate reliability and greater potential for post-refinement accuracy.
In our approach, methods incorporating DPO generally outperform SFT-only methods. Specifically, models trained with DPO Joint Training achieve higher Optimal-F1 scores, while those undergoing Post Training yield superior Quality-F1 scores. This suggests that Joint Training models adopt a more conservative stance, deferring to subsequent refinement, whereas Post Training models exhibit greater confidence, producing predictions of generally higher quality. This discrepancy stems from the differing sampling distributions of DPO data: Joint Training samples from a pre-SFT distribution, where performance is relatively poor, leading to a dataset with a higher proportion of ”unsure” predictions and resulting in more cautious model behavior.
Notably, some baselines, such as Self-Consistency, often achieve higher , , and Optimal-F1 scores compared to others, suggesting initial superiority. However, this indicates an undesirable scenario: an overly conservative model that frequently labels predictions as ”unsure.” This approach leads to a suboptimal balance between prediction types, evidenced by its significantly lower Quality-F1 score. In extreme cases, a model could assign ”unsure” to all predictions, thereby maximizing Optimal-F1.However, this would merely shift the burden to the refinement stage, reduce immediate usability and fails to balance reliability and helpfulness.
@@ 4.3 Refinement for Unsure Prediction
We trained an automatic refinement model for each dataset as described in Section 3.3. We refine the predictions labeled as ”unsure” from the initial stage and merge the refined outputs with the ”sure” predictions to generate the final results. The overall evaluation results are presented in Table 2. For RQ2, It can be observed that our method, after refinement, achieves significantly better performance with substantial improvements in Slot-F1 and JGA.
Method Type | Method | MultiWOZ-2.4 | BiTOD | SGD | |||
LLAMA3-8B | |||||||
Prompt | Direct | 78.86 | 36.01 | 82.40 | 36.46 | 32.41 | 13.58 |
Verbose | 76.80 | 36.69 | 76.93 | 27.99 | 23.11 | 9.00 | |
SR | 78.68 | 29.72 | 76.50 | 30.74 | 29.90 | 11.25 | |
Uncertainty | Prob | 75.89 | 24.19 | 73.88 | 24.86 | 28.48 | 9.63 |
SC | 70.98 | 20.70 | 69.98 | 24.86 | 25.60 | 4.39 | |
P(True) | 73.88 | 19.18 | 67.07 | 19.03 | 26.29 | 7.34 | |
Reliability (Ours) | SFT | 85.95 | 53.76 | 85.21 | 60.09 | 74.04 | 27.62 |
+ DPO Joint | 86.47 | 54.31 | 84.53 | 61.37 | 75.47 | 29.18 | |
+ DPO Post | 86.55 | 56.33 | 85.87 | 64.95 | 73.96 | 28.13 | |
Qwen2.5-7B | |||||||
Prompt | Direct | 78.66 | 36.35 | 83.23 | 37.39 | 42.02 | 16.93 |
Verbose | 67.59 | 26.45 | 74.37 | 27.38 | 27.79 | 9.19 | |
SR | 68.88 | 19.09 | 67.18 | 19.15 | 34.77 | 8.78 | |
Uncertainty | Prob | 74.18 | 23.56 | 73.97 | 25.64 | 35.91 | 10.81 |
SC | 59.34 | 23.45 | 65.60 | 28.31 | 34.34 | 3.06 | |
P(True) | 67.52 | 17.05 | 66.25 | 17.99 | 33.66 | 8.07 | |
Reliability (Ours) | SFT | 84.50 | 51.90 | 85.29 | 58.39 | 74.86 | 28.02 |
+ DPO Joint | 87.06 | 52.41 | 85.44 | 62.31 | 74.85 | 27.06 | |
+ DPO Post | 87.57 | 53.19 | 85.86 | 63.61 | 74.95 | 28.51 |
It is important to note that, referring to §4.2, despite a higher theoretical performance limit (Optimal-F1), DPO Joint Training models exhibit suboptimal Slot-F1 and JGA after refinement to Post Training models. This discrepancy stems from the limitations of the refinement model. Since the refinement model is not ideal, excessive ”unsure” predictions can lead to residual errors after refinement, incurring significant costs without proportional gains. This highlights the importance of a balanced ”sure” and ”unsure” classification. An ideal model should minimize unnecessary ”unsure” predictions and refine only when truly warranted, maximizing both efficiency and overall performance.
@@ 4.4 Scalability Analysis
To evaluate the scalability of our methods, we conducted experiments on foundation models with varying task capabilities. We consider the original LLAMA3 8B model a lower-performance baseline (denoted as Low) and applied supervised fine-tuning (SFT) on each dataset to enhance its DST capabilities. We train two foundation models using 1,000 samples (Medium) and 10,000 samples (High), then continually conduct reliability training and evaluate the performance of the obtained reliable LLM against baseline algorithms, with results shown in Figure 3. For RQ3, our methods consistently outperform baselines across different foundation models after refinement, demonstrating strong scalability—particularly in real-world scenarios where models are often domain-specific tuned and possess high task capabilities. Detailed results in Appendix §E further support these findings.
Additionally, as illustrated in Appendix Table 5, the advantage of DPO Post Training over Joint Training becomes more pronounced as foundation model performance improves. This trend stems from the offline nature of our DPO algorithm—higher-performing foundation models exacerbate the limitations of the Joint Training data distribution compared to Post Training, thereby affecting overall performance. Notably, in the SGD dataset, the benefits of DPO Joint Training diminish as the foundation model’s performance increases, even underperforming the SFT-only approach.
@@ 4.5 DPO Preference Strategies
We construct our DPO preference data using Weighted-F1 as the ranking metric, as detailed in §2.2.1. Weighted-F1 incorporates two adjustable parameters: influences Weighted Precision, while affects Weighted Recall. Using the Reliability-SFT trained LLAMA3 8B model as backbone, we further generated DPO preference data with varying values and conducted consecutive Post Training. Test results on the MultiWOZ-2.4 dataset (Figure 4) show that the model achieves optimal performance with and , reaching near-best results at a relatively low cost.
Analysis indicates that increasing heightens sensitivity to errors in ”unsure” predictions, reducing tolerance and yielding fewer ”unsure” outputs. Conversely, increasing amplifies rewards for successfully recalling ”unsure” predictions, encouraging the model to produce more ”unsure” predictions and adopt a more conservative stance. Practically, the optimal alpha settings depend on the scenario and refinement model performance. If the refinement model underperforms, excessive costs for marginal gains may not be justified. The flexible adjustment of both parameters allows for adaptation to diverse real-world scenarios.
@@ 4.6.1 Prediction Types Analysis
We further analyzed the model’s behavior after Reliability Training to evaluate its ability to make reliable confidence judgments in its outputs. Specifically, we examined the distribution of output categories and compared them to baseline models. These categories include five cases: correct and incorrect predictions classified as ”sure” or ”unsure”, along with missing predictions. The results are presented in Figure 5.
Our findings highlight two key improvements: (1) Enhanced reliability—The proportion of incorrect predictions within the ”sure” category has been notably reduced, with most errors now concentrated in the ”unsure” category. This indicates a stronger self-assessment capability, allowing the model to better identify potentially erroneous predictions compared to baselines. Additionally, across all scenarios, the ”unsure” category contains fewer correct predictions than incorrect ones, demonstrating the model’s effectiveness in placing accurate outputs in the ”sure” category, a distinction baseline models struggle to achieve. (2) Improved recall—Our approach significantly reduces missing predictions, capturing previously omitted relevant information within the ”unsure” category. While some may still contain errors, this shift enhances overall system helpfulness and provides a stronger foundation for further refinement.
@@ 4.6.2 Cost Analysis
In this work, ”Cost” refers to the additional overhead from the refinement process. We define the cost of refining an ”unsure” prediction as a constant value of 1, equating Cost to the number of ”unsure” predictions. For better comparison, we express Cost as the proportion of ”unsure” predictions.
We compare the performance and Cost of various baselines, as shown in Figure 6. Results are presented only for the MultiWOZ-2.4 dataset, with complete findings available in Appendix §F. Our methods demonstrate significant performance improvements while maintaining a relatively low Cost. Furthermore, as illustrated in Appendix Table 7, Cost decreases significantly with improved foundation model performance. This aligns with intuitive expectations: as the model strengthens, the proportion of uncertain predictions diminishes.
@@ 4.6.3 Refinement Model comparison
We employ various refinement models to enhance the predictions of the LLAMA3 Reliability-SFT+DPO Post model, comparing their performance as shown in Table 7. In addition to our trained refinement models, we utilize GPT-4o and the recently popular DeepSeek series (DeepSeek-AI et al., 2025), including the distilled 8B and 70B versions, all exhibiting strong reasoning capabilities. The results indicate significant performance variations among the different refinement models, with the DeepSeek 8B model underperforming, while the 70B model closely matches GPT-4o. Furthermore, our task-specific fine-tuning with CoT substantially improves refinement accuracy.
@@ 5 Related Work
Various methods have been proposed to assess these boundaries. Uncertainty-based approaches quantify prediction confidence through token probabilities (Manakul et al., 2023) and consistency (Chen & Mueller, 2024). Calibration strategies, including prompting (Tian et al., 2023) and fine-tuning (Tao et al., 2024) for improved confidence expression, align model confidence with prediction accuracy. Internal State Probing evaluates prediction factuality by analyzing model states, such as activations (Li et al., 2024a). While effective, these methods often rely on external post-hoc techniques, which are computationally expensive and lack integration with the model’s endogenous reasoning processes (Huang et al., 2024; Zhou et al., 2024).
To further align model behavior rather than merely estimate uncertainty, some studies guide models in self-reflection to minimize self-contradictions (Wei et al., 2022; Ji et al., 2023), albeit with increased computational overhead. More recently, RL-based approaches leverage uncertainty estimation to train reward models, enhancing LLM truthfulness by rejecting queries beyond their capabilities (Xu et al., 2024b; Chen et al., 2024; Xue et al., 2024). While effective in mitigating misinformation, these methods rely on the accuracy of uncertainty estimation algorithms and meanwhile may unintentionally reduce overall helpfulness.
@@ Appendix A Training Details
Appendix B Training Data Construction
@@
Input |
Dialog history: USER: Is there a modern European restaurant in the north? SYSTEM: I found city stop restaurant. They are located at Cambridge City Football Club Milton Road Chesterton. USER: That sounds lovely! SYSTEM: Would you like me to help you with reserving a table? USER: Maybe. Is that restaurant in the moderate price range? SYSTEM: No, it is in the expensive price range. Is this okay, or would you like to try another area? USER: How about a restaurant that serves Indian food? SYSTEM: There are several restaurants what part of town would you like it? USER: A restaurant in the north, please. Reliable dialog state: |
Gold Labels |
{’restaurant’: {’area’: ’north’, ’food’: ’indian’, ’pricerange’: ’moderate’}} |
LLM Original Predictions |
{’restaurant’: {’area’: ’north’, ’food’: ’european’, ’name’: ’Cambridge City Football Club Milton Road Chesterton’}} |
Constructed SFT Sample Output |
{’restaurant’: {’area’: {’value’: ’north’, ’confidence’: ’sure’}, ’food’: {’value’: ’European’, ’confidence’: ’unsure’}, ’pricerange’: {’value’: ’moderate’, ’confidence’: ’unsure’}, ’name’: {’value’: ’Cambridge City Football Club Milton Road Chesterton’, ’confidence’: ’unsure’}}} |
Explanation |
In the LLM’s original predictions, the value for area is correct, food is incorrect, name is extraneous, and pricerange is missing. Consequently, in the constructed SFT sample output, area is labeled as ”sure,” while the other three are marked as ”unsure,” with erroneous information left uncorrected. |
Input |
Dialog history: USER: Is there a modern European restaurant in the north? SYSTEM: I found city stop restaurant. They are located at Cambridge City Football Club Milton Road Chesterton. USER: That sounds lovely! SYSTEM: Would you like me to help you with reserving a table? USER: Maybe. Is that restaurant in the moderate price range? SYSTEM: No, it is in the expensive price range. Is this okay, or would you like to try another area? USER: How about a restaurant that serves Indian food? SYSTEM: There are several restaurants what part of town would you like it? USER: A restaurant in the north, please. Reliable dialog state: |
Gold Labels |
{’restaurant’: {’area’: ’north’, ’food’: ’indian’, ’pricerange’: ’moderate’}} |
Multi-Sampling With the Foundation LLM |
Constructed DPO Sample Output Pairs |
Chosen {’restaurant’: {’area’: {’value’: ’north’, ’confidence’: ’sure’}, ’food’: {’value’: ’indian’, ’confidence’: ’sure’}, ’pricerange’: {’value’: ’moderate’, ’confidence’: ’unsure’}}} Rejected {’restaurant’: {’area’: {’value’: ’north’, ’confidence’: ’sure’}, ’food’: {’value’: ’indian’, ’confidence’: ’sure’}}} |
Explanation |
Compared to the rejected sample, the chosen sample exhibits a higher recall for a correct pricerange within the ”unsure” category, aligning more closely with our alignment objectives. This is reflected in a higher Weighted-F1 score, justifying its selection over the rejected sample. |
Appendix C Refinement Model

### v4

@@ 4.2 Reliability Evaluation
Our approach consistently achieves superior Quality-F1 scores, demonstrating improved ability to distinguish predictions as “sure” or “unsure” based on confidence. A high indicates reliable “sure” predictions, fostering user trust with minimal need for verification. Simultaneously, a high ensures the “unsure” category contains a broad range of potentially correct outputs, which forms a solid basis for refinement. Together, these metrics lead to a higher Optimal-F1 score, representing the theoretical maximum post-refinement performance. In essence, our system optimally balances precision and recall, providing both immediate reliability and a strong potential for post-refinement accuracy.
In our approach, DPO-based methods generally outperform SFT-only methods. Specifically, models trained with DPO Joint Training achieve higher Optimal-F1 scores, suggesting a more conservative stance that defers to subsequent refinement. In contrast, those undergoing Post Training yield superior Quality-F1 scores, indicating greater confidence and higher-quality predictions. This discrepancy arises because Joint Training samples from a pre-SFT distribution, which has poorer performance. This results in a dataset with a higher proportion of “unsure” predictions, leading to more cautious model behavior.
Notably, some baselines, such as Self-Consistency, show high , , and Optimal-F1 scores, suggesting initial superiority. However, this often indicates an overly conservative model that frequently labels predictions as “unsure.” This leads to a suboptimal balance between prediction types, as evidenced by a significantly lower Quality-F1 score. In an extreme case, a model could label all predictions as “unsure”, maximizing Optimal-F1 but shifting the entire burden to the refinement stage. This approach sacrifices immediate usability and fails to balance reliability and helpfulness.
@@ 4.3 Refinement for Unsure Prediction
As described in Section §3.3, we trained an automatic refinement model for each dataset. We refined the “unsure” predictions from the initial stage and merged them with the “sure” predictions to generate the final results. Table 2 shows the overall evaluation. For RQ2, our method achieves significantly better performance after refinement, with substantial improvements in Slot-F1 and JGA.
MultiWOZ-2.4 | BiTOD | SGD | |||||
| Slot-F1 | JGA | Slot-F1 | JGA | Slot-F1 | JGA | ||||
| LLAMA3-8B | |||||||||
| Prompt | Direct | 78.86 | 36.01 | 82.40 | 36.46 | 32.41 | 13.58 | ||
| Verbose | 76.80 | 36.69 | 76.93 | 27.99 | 23.11 | 9.00 | |||
| SR | 78.68 | 29.72 | 76.50 | 30.74 | 29.90 | 11.25 | |||
| Uncertainty | Prob | 75.89 | 24.19 | 73.88 | 24.86 | 28.48 | 9.63 | ||
| SC | 70.98 | 20.70 | 69.98 | 24.86 | 25.60 | 4.39 | |||
| P(True) | 73.88 | 19.18 | 67.07 | 19.03 | 26.29 | 7.34 | |||
SFT | 85.95 | 53.76 | 85.21 | 60.09 | 74.04 | 27.62 | ||
| + DPO Joint | 86.47 | 54.31 | 84.53 | 61.37 | 75.47 | 29.18 | |||
| + DPO Post | 86.55 | 56.33 | 85.87 | 64.95 | 73.96 | 28.13 | |||
| Qwen2.5-7B | |||||||||
| Prompt | Direct | 78.66 | 36.35 | 83.23 | 37.39 | 42.02 | 16.93 | ||
| Verbose | 67.59 | 26.45 | 74.37 | 27.38 | 27.79 | 9.19 | |||
| SR | 68.88 | 19.09 | 67.18 | 19.15 | 34.77 | 8.78 | |||
| Uncertainty | Prob | 74.18 | 23.56 | 73.97 | 25.64 | 35.91 | 10.81 | ||
| SC | 59.34 | 23.45 | 65.60 | 28.31 | 34.34 | 3.06 | |||
| P(True) | 67.52 | 17.05 | 66.25 | 17.99 | 33.66 | 8.07 | |||
SFT | 84.50 | 51.90 | 85.29 | 58.39 | 74.86 | 28.02 | ||
| + DPO Joint | 87.06 | 52.41 | 85.44 | 62.31 | 74.85 | 27.06 | |||
| + DPO Post | 87.57 | 53.19 | 85.86 | 63.61 | 74.95 | 28.51 | |||
Referring to §4.2, despite their higher theoretical performance limit (Optimal-F1), DPO Joint Training models exhibit suboptimal Slot-F1 and JGA after refinement compared to Post Training models. This discrepancy stems from the limitations of the refinement model. Since the refinement model is not perfect, an excessive number of “unsure” predictions can lead to residual errors after refinement, incurring significant costs without proportional gains. This highlights the importance of a balanced “sure” and “unsure” classification. An ideal model should minimize unnecessary “unsure” predictions and refine only when truly warranted, maximizing both efficiency and overall performance.
@@ 4.4 Scalability Analysis
To evaluate scalability, we experimented with foundation models of varying capabilities. We consider the original LLAMA3 8B as a low-performance baseline (denoted as Low), then fine-tuned two models with 1,000 (Medium) and 10,000 (High) samples to enhance their DST capabilities. We then continually conducted reliability training and evaluated the performance of the obtained reliable LLMs against baseline methods. As shown in Figure 3, our methods consistently outperform baselines after refinement, demonstrating strong scalability. This is particularly relevant for real-world scenarios where models are often domain-specifically tuned and possess high task capabilities. Detailed results in Appendix §F further support these findings.
As illustrated in Table 7 in Appendix §F, the advantage of DPO Post Training over Joint Training becomes more pronounced as foundation model performance improves. This is because our offline DPO algorithm’s limitations are exacerbated by higher-performing foundation models, negatively affecting the Joint Training data distribution and overall performance. Notably, in the SGD dataset, the benefits of DPO Joint Training diminish with increasing foundation model performance, even falling behind the SFT-only approach.
@@ 4.5 DPO Preference Strategies
We constructed our DPO preference data using Weighted-F1 as the ranking metric, as detailed in §2.2.1. Weighted-F1 incorporates two adjustable parameters: for Weighted Precision and for Weighted Recall. Using the Reliability-SFT trained LLAMA3 8B model, we generated DPO preference data with varying values and performed subsequent Post Training. Results on the MultiWOZ-2.4 dataset (Figure 4) show that the model achieves optimal performance at and , reaching near-best results at a relatively low cost.
Analysis indicates that increasing heightens sensitivity to errors in “unsure” predictions, reducing tolerance and the number of “unsure” outputs. Conversely, increasing amplifies rewards for successfully recalling “unsure” predictions, encouraging a more conservative stance. The optimal alpha settings depend on the specific scenario and refinement model performance; if the refinement model underperforms, the cost of excessive “unsure” predictions may not be justified. The flexible adjustment of both parameters allows for adaptation to diverse real-world scenarios.
@@ 4.6.1 Prediction Types Analysis
To evaluate our model’s ability to make reliable confidence judgments, we further analyzed the distribution of its output categories and compared them to baselines. We examined five output categories: correct and incorrect predictions classified as “sure” or “unsure”, along with missing predictions. The results are presented in Figure 5.
Our findings highlight two key improvements: (1) Enhanced reliability: We observed a notable reduction in incorrect “sure” predictions, with most errors concentrated in the “unsure” category. This indicates a stronger self-assessment capability, allowing the model to better identify potential errors than baselines. Additionally, the “unsure” category consistently contains fewer correct predictions than incorrect ones, a distinction baseline models struggle to achieve. (2) Improved recall: Our approach significantly reduces missing predictions by capturing previously omitted relevant information within the “unsure” category. While some of these may still contain errors, this shift enhances overall helpfulness and provides a stronger foundation for refinement.
@@ 4.6.2 Cost Analysis
In this work, Cost refers to the additional overhead from the refinement process. We define the cost of refining an “unsure” prediction as a constant value of 1, equating Cost to the number of “unsure” predictions. For better comparison, we express Cost as the proportion of “unsure” predictions.
We compare the performance and Cost of various baselines in Figure 6. Results are shown for the MultiWOZ-2.4 dataset, with complete findings in Appendix §G. Our methods show significant performance improvements at a relatively low Cost. Additionally, as shown in Appendix Table 9, Cost decreases significantly as the foundation model’s performance improves, which is expected as a stronger model produces fewer uncertain predictions.
@@ 4.6.3 Refinement Model Comparison
We used various refinement models to enhance predictions from the LLAMA3 Reliability-SFT+DPO Post model. As shown in Table 7, we compared our trained models against GPT-4o and the DeepSeek series (DeepSeek-AI et al., 2025) (8B and 70B), all known for strong reasoning. Results show significant performance differences, with DeepSeek 8B underperforming while the 70B model nearly matches GPT-4o. Our task-specific fine-tuning with CoT substantially improves refinement accuracy.
4.7 Extension to Mathematical Problems
To validate the generalization of our method to different tasks, we conducted supplementary experiments on the widely used GSM8K mathematics dataset (Cobbe et al., 2021) using the Qwen2.5-7B-Instruct model with 1,000 training samples. We evaluated the model under two test settings: (a) with Chain-of-Thought (CoT) reasoning (Wei et al., 2023) and (b) by directly predicting the final answer. We tested both settings as LLMs generally perform differently on each. For the refinement model, we used the DeepSeek-R1-Distill-Llama-70B model, which has strong mathematical reasoning capabilities.
We adapted the F1-based evaluation metric from Section §2.2 to an accuracy-based definition:
As shown in Table 3, our method consistently outperformed both prompt-based and uncertainty-based baselines in improving the model’s self-assessment ability, achieving higher Quality-Accuracy. Furthermore, after applying refinement with the same refinement model, our approach yielded a higher final accuracy at a reasonable computational cost.
| Method Type | Method | Quality Accuracy | Final Accuracy | ||
| Prompt | Direct | 86.81 | 86.81 | ||
| Verbose | 90.8 | 90.98 | |||
| Uncertainty | Prob | 87.69 | 88.55 | ||
| SC | 89.41 | 91.48 | |||
SFT | 89.69 | 91.05 | ||
| + DPO Joint | 90.14 | 93.10 | |||
| + DPO Post | 92.73 | 95.30 |
| Method Type | Method | Quality Accuracy | Final Accuracy | ||
| Prompt | Direct | 19.94 | 19.94 | ||
| Verbose | 21.44 | 24.34 | |||
| Uncertainty | Prob | 24.49 | 76.50 | ||
| SC | 24.37 | 70.05 | |||
SFT | 27.31 | 77.94 | ||
| + DPO Joint | 26.16 | 78.24 | |||
| + DPO Post | 27.83 | 79.83 |
@@ 5 Related Work
Various methods have been proposed to assess these boundaries. Uncertainty-based approaches quantify prediction confidence through token probabilities (Manakul et al., 2023) and consistency (Chen & Mueller, 2024). Calibration strategies, including prompting (Tian et al., 2023) and fine-tuning (Tao et al., 2024) for improved confidence expression, align model confidence with prediction accuracy. Internal State Probing evaluates prediction factuality by analyzing model states, such as activations (Li et al., 2024a). While effective, these methods often rely on external post-hoc techniques, which are computationally expensive and lack integration with the model’s endogenous reasoning processes (Huang et al., 2024; Zhou et al., 2024; Pan et al., 2025; Shan et al., 2025).
To further align model behavior rather than merely estimate uncertainty, some studies guide models in self-reflection to minimize self-contradictions (Wei et al., 2022; Ji et al., 2023b), albeit with increased computational overhead. More recently, RL-based approaches leverage uncertainty estimation to train reward models, enhancing LLM truthfulness by rejecting queries beyond their capabilities (Xu et al., 2024b; Chen et al., 2024; Xue et al., 2024). While effective, these methods rely on the accuracy of uncertainty estimation algorithms and meanwhile may unintentionally reduce overall helpfulness.
@@ 6 Conclusion
Acknowledgments
This work is funded by the China NSFC Projects (62120106006, 92370206, and U23B2057) and Shanghai Municipal Science and Technology Projects (2021SHZDZX0102 and 25X010202846).
@@ Appendix A Training Details
Appendix B Execution Example for Two Stages
We provide an example sample that includes the complete execution flow and model outputs for the two stages as shown in Table 4.
| Input | ||||||||
| Stage One: Reliable LLM Predictions | ||||||||
| Stage Two: Refinement for “Unsure” Predictions: | ||||||||
| Final Refined Predictions: | ||||||||
Appendix C Training Data Construction
@@
| Input | |||||||||||
| Gold Labels | |||||||||||
| LLM Original Predictions | |||||||||||
| Constructed SFT Sample Output | |||||||||||
| Explanation | |||||||||||
| Input | |||||||||||
| Gold Labels | |||||||||||
| Multi-Sampling With the Foundation LLM | |||||||||||
| Constructed DPO Sample Output Pairs | |||||||||||
| Explanation | |||||||||||
Appendix D Refinement Model